# XRISM Observations of The Prototypical Cold Front in Abell 3667

### v2

@@ 2.1 Observations
XRISM/Resolve conducted two observations of A3667 from 30 October to 4 November 2025. The Resolve field of view (FoV) is a square covered by an array of pixels, each one returning a spectrum with an average resolution of 4.5 eV (FWHM), (M. Tashiro et al., 2025). The two observed fields are indicated by white boxes in Figure 1. One pointing covered the edge of the cold front (CF_IN; OBSID: 201050010, RA, Dec = 303.26525, -56.88616; PI: Ichinohe), and the other targeted the cluster center (Center; OBSIDs: 201051010, RA, Dec = 303.17404, -56.84815; PI: Ichinohe).
@@ 2.2 Data reduction and Spectral modeling
Data reduction was performed following the procedure in Y. Omiya et al. (2025). We processed the data with xapipeline in HEAsoft v6.34 (CALDB v11), applied the screening criteria, and generated RMFs with rslmkrmf and ARFs with xaarfgen using the Chandra image in the keV band. After screening, the net exposures were 105 ks for Center and 276 ks for CF_IN. Instrumental systematics were estimated using the in-orbit 55Fe calibration data, yielding uncertainties of 15 km s-1 for line positions and km s-1 for line broadening at the Fe K line energies, both negligible relative to the statistical errors.
Spectral fitting was performed in XSPEC v12.15.0 (K. A. Arnaud, 1996) with AtomDB 3.0.9, adopting protosolar abundances from K. Lodders et al. (2009). The ICM emission was modeled with bapec under collisional ionization equilibrium, with free parameters of temperature, metal abundance, redshift, and Gaussian velocity dispersion. Average heliocentric velocity corrections of km s-1, derived using barycen, were applied to the fits for both pointings. Galactic absorption was modeled with tbabs with a fixed column density of cm-2 (P. M. W. Kalberla et al., 2005). Non–X-ray background (NXB) spectra were generated with rslnxbgen from night-Earth events using the Resolve NXB Database (v2), modeled with a power-law and Gaussian lines, and the best-fit model was included as the NXB component in the fits.
@@ 3.2 Resolving the cluster velocity structure
The sky regions were chosen to probe interesting physical features in the cluster. Figure 1 shows a Chandra broad-band image and a soft-to-hard ratio (SHR) map (a ratio of the 0.5–2.0 keV and 2.0–7.0 keV images), which shows variations of the projected ICM temperature. The sky regions are indicated by the orange dashed contours in Figure 1. The region outside (on the less-dense side of) the cold front is defined as region A. Immediately inside the front, we identify a bright cool region, designated as region B. Moving further northwest from the cold front along the cluser long axis, the SHR map reveals two step-like temperature changes (also seen in the temperature map in Y. Ichinohe et al. 2017). Regions C–G were defined to trace those steps and also probe the structure across the cluster’s long axis.
@@
The spectrum of the detector region a, which covers part of the sky region A, suffers from limited photon statistics and a significant scattered contribution from the adjacent, brighter region B. The temperature determination based solely on the continuum is susceptible to systematic error. We therefore fixed the temperature of region A at 8.0 keV and the abundance at 0.41 , based on results of Y. Omiya et al. (2024). As discussed in Appendix A, this does not alter our main conclusions.
@@ Acknowledgements
Y.O. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. Y.O. was supported by the Sasakawa Scientific Research Grant from The Japan Science Society. This work was supported by JSPS KAKENHI grant numbers JP20H00157 (K.N.), JP22H00158, JP23H04899, JP25H00672 (Y.F.), and JP25K23398 (S.U.). The material is based upon work supported by NASA under award number 80GSFC21M0002 (F.M.). S.U. acknowledges support by Program for Forming Japan’s Peak Research Universities (J-PEAKS) Grant Number JPJS00420230006. I.Z. acknowledges support from NASA grant 80NSSC18K1684 and partial support from the Alfred P. Sloan Foundation through the Sloan Research Fellowship. Support for J.Z. was provided by the Chandra X-ray Observatory Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060.
@@ Appendix A Robustness of the SSM analysis in region A
Figure 5 shows the resulting maps. In region A, the free fit yields a temperature of keV and an abundance of , indicating a cooler and more metal-rich solution compared to the fixed-parameter analysis. However, the bulk velocity and velocity dispersion remain consistent within statistical uncertainties, and the overall velocity structure is unchanged. This can be clearly seen in Figure 2, where the spectrum of (detector plane) region a, excluding the region A model (i.e., the sum of all other models), exhibits positive residuals on the high-energy side of the He-like Fe line complex. These residuals require the inclusion of a blueshifted component from region A, demonstrating that the velocity offset is robust.

### v3

@@ 2.1 Observations
XRISM/Resolve conducted two observations of A3667 from 30 October to 4 November 2024. The Resolve field of view (FoV) is a square covered by an array of pixels, each one returning a spectrum with an average resolution of 4.5 eV (FWHM), (M. Tashiro et al., 2025). The two observed fields are indicated by white boxes in Figure 1. One pointing covered the edge of the cold front (CF_IN; OBSID: 201050010, RA, Dec = 303.26525, -56.88616; PI: Ichinohe), and the other targeted the cluster center (Center; OBSIDs: 201051010, RA, Dec = 303.17404, -56.84815; PI: Ichinohe).
@@ 2.2 Data reduction and Spectral modeling
Data reduction was performed following the procedure in Y. Omiya et al. (2025) using XRISM CALDB version 20250315. We processed the data with xapipeline in HEAsoft v6.34 (CALDB v11), applied the screening criteria, and generated RMFs with rslmkrmf and ARFs with xaarfgen using the Chandra image in the keV band. After screening, the net exposures were 105 ks for Center and 276 ks for CF_IN. Instrumental systematics were estimated using the in-orbit 55Fe calibration data, yielding uncertainties of 15 km s-1 for line positions and km s-1 for line broadening at the Fe K line energies, both negligible relative to the statistical errors.
Spectral fitting was performed in XSPEC v12.15.0 (K. A. Arnaud, 1996) with AtomDB 3.0.9, adopting protosolar abundances from K. Lodders et al. (2009). The ICM emission was modeled with bapec under collisional ionization equilibrium, with free parameters of temperature, metal abundance, redshift, and Gaussian velocity dispersion. Average heliocentric velocity corrections of km s-1, derived using barycen, were applied to the fits for both pointings. Galactic absorption was modeled with tbabs with a fixed column density of cm-2 (P. M. W. Kalberla et al., 2005). Non–X-ray background (NXB) spectra were generated with rslnxbgen from night-Earth events using the Resolve NXB Database (v2). They were modeled with a power-law and Gaussian lines, and the best-fit model was included as the NXB component in the fits, following the recommended procedures described in the “Resolve NXB Database and Spectral Extraction Recipes” on the HEASARC website.
@@ 3.2 Resolving the cluster velocity structure
The sky regions were chosen to probe interesting physical features in the cluster. Figure 1 shows a Chandra broad-band image and a soft-to-hard ratio (SHR) map (a ratio of the 0.5–2.0 keV and 2.0–7.0 keV images), which shows variations of the projected ICM temperature. The sky regions are indicated by the orange dashed contours in Figure 1. The region outside (on the less-dense side of) the cold front is defined as region A. Immediately inside the front, we identify a bright cool region, designated as region B. Moving further northwest from the cold front along the cluser long axis, the SHR map reveals two step-like temperature changes (also seen in the temperature map in Y. Ichinohe et al. 2017). Regions C–G were defined to trace the successive steps in the SHR map, from the immediate vicinity of the cold front to the downstream outskirts, as well as to probe the thermodynamic and kinematic variations along the cluster’s long (merger) axis. Specifically, Regions C and D lie just inside the cold front and are separated into the right (north) and left (south) sides relative to the inferred subcluster trajectory. Regions E and F are located at intermediate distances and are similarly divided into two sides, while Region G covers the outermost downstream sector.
@@
The spectrum of the detector region a, which covers part of the sky region A, suffers from limited photon statistics and a significant scattered contribution from the adjacent, brighter region B. The temperature determination based solely on the continuum is susceptible to systematic error. We therefore fixed the temperature of region A at 8.0 keV and the abundance at 0.41 , based on results of Y. Omiya et al. (2024). As discussed in Appendix 5, this does not alter our main conclusions.
@@ 4.4 Constraints on magnetic field strength
We emphasize that this estimate assumes, for consistency with the merger geometry adopted throughout this work, that the velocity contrast arises purely from motions along the line of sight, with no sky–plane component of the shear. Under this simplified assumption, the derived should be interpreted as a rough estimate, because any unmodeled geometrical component of the flow could modify the true tangential shear velocity.
@@ Acknowledgements
Y.O. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. Y.O. was supported by the Sasakawa Scientific Research Grant from The Japan Science Society. This work was supported by JSPS KAKENHI grant numbers JP20H00157 (K.N.), JP22H00158, JP23H04899, JP25H00672 (Y.F.), and JP25K23398 (S.U.). The material is based upon work supported by NASA under award number 80GSFC21M0002 (F.M.). S.U. acknowledges support by Program for Forming Japan’s Peak Research Universities (J-PEAKS) Grant Number JPJS00420230006. I.Z. acknowledges support from NASA grant 80NSSC18K1684 and partial support from the Alfred P. Sloan Foundation through the Sloan Research Fellowship. Support for J.Z. was provided by the Chandra X-ray Observatory Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection https://doi.org/10.25574/cdc.503 (catalog DOI: 10.25574/cdc.503)
@@ Appendix A Robustness of the SSM analysis in region A
Figure 5 shows the resulting maps. In region A, the free fit yields a temperature of keV and an abundance of , indicating a cooler and more metal-rich solution compared to the fixed-parameter analysis. As shown in Table 2, the simple single–temperature fits that ignore SSM mixing yield notably different thermodynamic parameters; for example, region a is fitted with a temperature of 4.5 keV, substantially lower than the value obtained from the free-parameter SSM analysis. This demonstrates that PSF and spatial mixing can bias the apparent temperature structure if not modeled properly. However, the bulk velocity and velocity dispersion remain consistent within statistical uncertainties, and the overall velocity structure is unchanged. This can be clearly seen in Figure 6, where the spectrum of (detector plane) region a, excluding the region A model (i.e., the sum of all other models), exhibits positive residuals on the high-energy side of the He-like Fe line complex. These residuals require the inclusion of a blueshifted component from region A, demonstrating that the velocity offset is robust.
@@
| Parameters | Regions a | b | c | d | e | f | g |
| Abundance () | |||||||
| Redshift (10-2) | |||||||
| Relative Velocity (km s-1) | |||||||
| Velocity dispersion (km s-1) | |||||||
| C-stat/ |